log base 2 (log base 3 (log base 5 of x)) = 0
Answer by stanbon(75887)

log base 2 (log base 3 (log base 5 of x)) = 0
-----
log2[log3(log(x)/log(5)) = 0
-----
log2[log[log(x)/log(5)]/log(3)] = 0
-----
log[log[log[log(x)/(log(5)*log(3))]/(log(2)) = 0
-----
log[log[log(x)/(log(5)*log(3)*log(2))] = 0
-----
log[log(x)/(log(5)*log(3)*log(2))] = 1
log(x)/(log(5)(log(3)*log(2))/ = 10
log(x) = 10*log(5)*log(3)(log(2))
log(x) = 1.0039
x = 10*1.0039
x = 10.0906
==============
Cheers,
Stan H.
-------------------